# Privacy Preservation Techniques (PPTs) in IoT Systems: A Scoping Review and Future Directions Author’s draft for soliciting feedback - February 20, 2025

### v1

@@ Abstract
Privacy preservation in Internet of Things (IoT) systems requires the use of privacy-enhancing technologies (PETs) built from innovative technologies such as cryptography and artificial intelligence (AI) to create techniques called privacy preservation techniques (PPTs). These PPTs achieve various privacy goals and address different privacy concerns by mitigating potential privacy threats within IoT systems. This study carried out a scoping review of different types of PPTs used in previous research works on IoT systems between 2010 and early 2023 to further explore the advantages of privacy preservation in these systems. This scoping review looks at privacy goals, possible technologies used for building PET, the integration of PPTs into the computing layer of the IoT architecture, different IoT applications in which PPTs are deployed, and the different privacy types addressed by these techniques within IoT systems. Key findings, such as the prominent privacy goal and privacy type in IoT, are discussed in this survey, along with identified research gaps that could inform future endeavors in privacy research and benefit the privacy research community and other stakeholders in IoT systems.
Index Terms:
@@
I Introduction
As the Internet of Things (IoT) continues to integrate into human daily life, effective privacy preservation techniques (PPTs) have become essential [1]. These PPTs mitigate privacy risks by protecting sensitive personal information within IoT systems.
The Internet of Things (IoT) concept dates back to the 1970s, when it evolved from a closed proprietary telemetry system, as explained by McKelvey et al. [2]. However, significant attention to privacy preservation within IoT began to grow after 1999, when the term ”Internet of Things” was first coined by Kevin Ashton [2].
@@ I Introduction
This increase in privacy focus is evident in the growing number of related publications on privacy in general and IoT systems from 1960 to early 2023, as shown in Figure 1. Figure 1 was generated from an analysis of privacy-related publications collected from Google Scholar using Harzing’s Publish or Perish software, designed to retrieve and analyze academic citations[3].
@@
Figure 1: The line graph shows publication trends for Privacy Preservation Techniques. The blue line indicates general privacy publications, while the green line shows privacy-related publications in IoT (Internet of Things) systems.
@@
From the analysis in this study, trends in IoT-related privacy publications started in 2002, as seen in Figure 1, with a significant increase from 2008 to 2023. This increase highlights increased awareness and recognition of the privacy risks associated with IoT systems.
It is important to note that, as illustrated in Figure 1, the collection of data was concluded in 2023. This excluded any recent publications from the second quarter of 2023 until now, as evidenced by the declining slope at the end of the line graph.
@@
As mentioned in some literature, technological advancements contribute to the new privacy challenges by exacerbating privacy risks [4]. Threat actors can now launch a privacy threat leveraging these new technologies such as Generative AI [4, 5]. However, technological advancements have also facilitated the development of sophisticated PPTs [6, 7]. Some of these technologies, such as artificial intelligence (AI), post-quantum cryptography, and blockchain, have been leveraged to build privacy-enhancing technologies (PETs) used to create PPTs that mitigate privacy threats.
@@
This study leverages scoping review, a systematic method for mapping existing evidence on a topic, by identifying key concepts, theories, and knowledge gaps [8]. This study embraces this approach due to its comprehensive overview of a subject area, helping researchers and practitioners to understand current knowledge and highlighting areas for further investigation.
@@
This study aims to use a scoping review to analyze various PPTs in the context of IoT, focusing on the technologies used to build PET, PPT implementation in the computing layers of IoT architecture, privacy protection types, different IoT applications in which PPTs are implemented, and privacy goals of these PPTs. This study addresses gaps not covered by previous reviews on PPT in IoT systems.
Contributions This review will focus on the PPTs related to IoT from the literature. At the time this study was conducted, and to the best of our knowledge, we were unaware of any publication explicitly focusing on a scoping review of PPTs used in IoT applications. Our contribution will focus on
@@
  • providing a scoping review of state-of-the-art PPTs in IoT systems from paper publications in the past thirteen and a half (13.5) years.
@@
  • comparison of each PPT from the literature to identify the most frequently used and least used PPTs across different IoT applications.
@@
This paper is organized as follows: Section I explains the trends in general privacy publications and those specific to privacy in IoT systems. Section II presents a brief background on privacy and key privacy terms in IoT systems, Section III shows the uniqueness of this study by comparing it to previous reviews on PPTs. Section IV explains the scoping review guideline and framework used in this study. Section V identifies the research questions in this study, while Section VI explains how we gathered relevant studies in this scoping review. Section VII presents the databases and the year range of selection of papers reviewed in this study. Section VIII categorizes the collected papers into five key areas our review focused on: PPT goals, technologies used to build PET, IoT applications and their PPTs, IoT computing layer in which PPTs are implemented, and privacy types which these PPTs. Section IX summarizes our findings, providing answers to the research questions in our study. Section X discusses the future direction based on the research questions in our study. Lastly, Section XI is the conclusion of the study, summarizing the outcome of the study and the limitations of this study.
II Definitions of key concepts
II-A Privacy
Privacy has evolved significantly since its origins in the 14th through 18th centuries. Initially focused on protecting one’s body and home, it transitioned by the end of the 19th century to include the control and protection of personal information [9]. Privacy refers to an individual’s ability to determine how much of their thoughts and actions they share with others and under what circumstances. This concept is closely tied to personal identity and is viewed as safeguarding an individual’s autonomy and independence [10].
II-B Types of Privacy Protection
@@
Privacy in the IoT revolves around users and other entities interacting within IoT systems. Finn et al. [11] identify seven types of privacy in technology. These types include the privacy of the person, the privacy of personal behavior, the privacy of personal communication, the privacy of personal data, the privacy of location and space, the privacy of thoughts and feelings, and the privacy of associations. However, not all of these privacy types are relevant to IoT. Therefore, based on existing literature, this study outlined the specific types of privacy in IoT systems.
@@
  • Device Identity privacy: To prevent unauthorized device tracking and smart device information disclosure such as firmware version, operating system, MAC address, etc. [12].
@@
  • Location privacy: To protect the user and smart device location information. This could prevent threats such as profiling [11, 12].
@@
  • Group privacy: To protect the privacy of a group of IoT system users [11]. For example, a family member of a Smart Home (SH) owner or department member in an Industrial IoT (IIoT) that has access to operate an automated system in an industry.
II-C Privacy Threats in IoT
@@
Depending on the threat actor, some privacy threats can be malicious, while some can be non-malicious. Privacy threats such as the identification of the subject, linkage, data leakage, inventory attacks, eavesdropping, life cycle transition, privacy violation during the presentation, and data tampering could be malicious due to the threat actor involved [13, 14, 15, 16]. In contrast, privacy threats such as Localization and tracking, and profiling are not usually malicious when they involve actors such as service providers, security agents, and third-party providers [13, 14].
II-D Privacy Preservation Techniques (PPT)
@@
PPTs are methods or processes implemented in IoT to achieve privacy goals [17]. These privacy goals vary depending on the techniques and the IoT application involved. These privacy goals include data masking, pseudonymization, data perturbation, multipath routing, and access controls, as shown in Table III. To achieve this goal by PPTs, the right technological building block needs to be developed, which are referred to as privacy enhancing technologies (PET) [18, 19, 20, 21]. PET in IoT may be challenging due to IoT systems’ constraints, such as memory capacity, processing power, etc. However, implementing these PETs in a lightweight manner will make them suitable for IoT systems [17].
II-E Privacy Enhancing Technologies (PET)
@@
These are technologies used to achieve PPTs in an IoT system. In IoT, PPTs have evolved, employing diverse PETs (such as differential privacy, Homomorphic encryption, pseudonymization, and secure multi-party computation, to mention a few) to provide different privacy preservation solutions [22, 23, 24, 25, 26, 27, 28].
II-F General Technology used in achieving Privacy Preservation Approach in IoT
@@
The Internet of Things (IoT) utilizes various privacy preservation approaches to achieve specific privacy goals and types. These approaches often leverage PETs developed through technological methods like blockchain [29], machine learning/deep learning [30], cryptography [31], informed consent [32], quantum-safe algorithms [33], mathematical models [34], and privacy-based infrastructure.
III Comparison with existing surveys and reviews
This section compares previous related surveys, emphasizing key parameters such as the technology used to build PET, goals of PPT, specific IoT applications, and the implementation of privacy and PPT across various IoT computing layers.
It is important to provide a brief background on the types of literature reviews and which literature review is used in our study.
Types of literature review: There are 14 types of literature reviews as explained by Grant et al. [35], which can be categorized into four main types: narrative, rapid, scoping, and systematic. A narrative review covers a small selected part of a study to support empirical research. However, replicating it is often challenging and may yield biased results due to its non-comprehensive nature[35, 36]. Rapid reviews quickly summarize and synthesize research findings within the constraints of time and resources, systematically striving for comprehensiveness and details [35, 37]. Systematic reviews aim to identify, appraise, and synthesize all empirical evidence that meets specific inclusion criteria to address focused research questions [35, 36]. Scoping reviews employ a comprehensive search strategy to map the literature within a broad context but do not synthesize evidence, often serving as a preliminary step before conducting a more focused systematic review [35, 38]. This study employs a scoping review methodology to map and analyze existing literature on PPTs within IoT systems.

### v2

@@ Abstract
Privacy preservation in Internet of Things (IoT) systems requires the use of privacy-enhancing technologies (PETs) built from innovative technologies such as cryptography and artificial intelligence (AI) to create techniques called privacy preservation techniques (PPTs). These PPTs achieve various privacy goals and address different privacy concerns by mitigating potential privacy threats within IoT systems. This study carried out a scoping review of different types of PPTs used in previous research works on IoT systems between 2010 and 2025 to further explore the advantages of PPT methods in these systems. This scoping review looks at PPT methods to achieve PPT goals, possible technologies used for building PET, the integration of PPTs into the computing layer of the IoT architecture, different IoT applications in which PPTs are deployed, and the different privacy types addressed by these techniques within IoT systems. Key findings, such as the prominent privacy goal and privacy type in IoT, are discussed in this survey, along with identified research gaps that could inform future endeavors in privacy research and benefit the privacy research community and other stakeholders in IoT systems.
keywords:
@@
journal: Computer & Security
[…]
organization=School of Information Technology,addressline=Carleton University, city=Ottawa, postcode=K1S 5B6, state=Ontario, country=Canada
1 Introduction
@@
This increase in privacy awareness is evident in the growing number of publications related to general privacy and those specifically focused on IoT systems from 1960 to early 2025, as shown in Figure 1. Figure 1 was generated from an analysis of privacy-related publications collected from Google Scholar using Harzing’s Publish or Perish software, designed to retrieve and analyze academic citations [190].
@@
Figure 1: The line graph shows publication trends for PPTs. The blue line indicates the publications on general privacy, while the green line shows privacy-related publications in IoT systems.
@@
According to the analysis in this study, trends in IoT-related privacy publications began in 2002, as shown in Figure 1, with a significant increase from 2008 to 2025. This increase highlights increased awareness and recognition of the privacy risks associated with IoT systems.
@@
As mentioned in some literature, technological advancements contribute to the new privacy challenges by exacerbating privacy risks [184]. Threat actors can now launch a privacy threat leveraging these new technologies, such as Generative AI [184, 99]. However, technological advancements have also facilitated the development of sophisticated PPTs [110, 323]. Some of these technologies, such as artificial intelligence (AI) and post-quantum cryptography, have been leveraged to build privacy-enhancing technologies (PETs) that are used to create PPTs, which mitigate privacy threats. A proper understanding of the PPTs development, application approach, and where it has been applied in IoT is essential to enlighten researchers on areas that require further research.
@@
This insight motivates our study by leveraging a scoping review, a systematic method for mapping existing evidence on a topic, which identifies key concepts, theories, and knowledge gaps [510]. This study adopts this approach due to its comprehensive overview of a subject area, enabling researchers and practitioners to understand the current state of knowledge and identify areas for further investigation.
@@
Our aim is to use a scoping review to analyze various PPTs in the context of IoT, focusing on the technologies used to build PET, PPT implementation in the computing layers of IoT architecture, privacy protection types, different IoT applications in which PPTs are implemented, and privacy goals of these PPTs. This study addresses gaps not covered by previous reviews on PPT in IoT systems. The contributions of this study is summarized as follows:
@@
  • A scoping review of state-of-the-art PPTs in IoT systems from paper publications in the past fourteen (14) years.
@@
  • A comparison of each PPT from the literature to identify the most frequently used and least used PPTs across different IoT applications.
@@
This paper is organized as follows: Section 1 explains the trends in general privacy publications and those specific to privacy in IoT systems. Section 2 presents a brief background on privacy and key privacy terms in IoT systems, Section 3 shows the uniqueness of this study by comparing it to previous reviews on PPTs. Section 4 explains the scoping review guideline and framework used in this study. Section 5 identifies the research questions in this study, while Section 6 explains how we gathered relevant studies in this scoping review. Section 7 presents the databases and the year range of selection of papers reviewed in this study. Section 8 categorizes the collected papers into five key areas our review focused on: PPT goals, technologies used to build PET, IoT applications and their PPTs, IoT computing layer in which PPTs are implemented, and privacy types in which these PPTs operate. Section 9 summarizes our findings and provides answers to the research questions in our study. Section 10 discusses the future direction based on the research questions in our study. Lastly, Section 11 presents the study’s conclusions and its limitations.
2 Overview of privacy in IoT systems
This section explores the fundamental concepts underlying privacy preservation within the Internet of Things (IoT) context. We begin by defining the concept of IoT and its architecture and elucidating the essence of communication modes in IoT. To fully appreciate the significance of PPTs in IoT, it is essential to establish a comprehensive understanding of privacy within the context of these techniques in IoT. Lastly, we examine the interplay between privacy terms such as PPT, PET, privacy types, privacy goals, and privacy threats in IoT systems.
This section will offer insights into the evolving landscape of PPTs in IoT, shedding light on the ever-changing dynamics of privacy and its safeguarding mechanisms.
2.1 Internet of Things (IoT)
IoT consists of ”things” that can interact with each other by communicating and exchanging data and information based on the senses about the environment that trigger actions in the form of services with little or no human intervention [487].
These ”things” are connected sensors and actuating devices that form a flexible network that effortlessly shares information across different platforms. This system operates within a unified framework, encouraging the development of a shared operational perspective and driving the creation of advanced applications [178]. This progress is the result of combining ubiquitous sensing, data analytics, and information representation, with cloud computing being the central and cohesive foundation for this integrated ecosystem [178]. IoT was first presented in 1999 by Kevin Ashton, and with technology development such as AI, it has evolved into a platform that influences our day-to-day activities. Its application has spanned various areas, including medicine, industry, agriculture, education, and energy, making operations in these areas more intelligent [66]. IoT systems comprise connecting IoT devices and smart devices, and are implemented using various architectures that cater to the specific needs of IoT applications.
2.2 IoT Devices and Smart Devices
IoT devices are physical objects in an IoT system that are connected to each other and to the internet to collect, send, receive, and store data. An example of an IoT device is a smart TV, a smart watch, a smartphone, to name a few. However, smart devices are intelligent objects, also known as smart objects, that possess unique identities, sensing and storage capabilities, data availability, communication abilities, and decision-making capabilities [301]. Smart devices are a subset of IoT devices and do not require human interaction. An example is a weather probe that collects weather data, and an IoT security system that involves a surveillance camera that communicates with other devices and sends alerts and footage in case of intrusion, leveraging Artificial Intelligence (AI) for asset tracking and monitoring, as seen in Industrial IoT (IIoT) [142]. Although the data from these smart devices is eventually collected by humans for analysis. The analyzed data can be categorized as either user or device data.
2.3 User and Device data in IoT System
Various data in the IoT system are processed and stored, and then exchanged between IoT devices during their interactions. These data can be categorized as either user data or device data.
  • User data: Data related to IoT users, which may be sensitive (e.g., personal identifiable information [PII]) or non-sensitive (e.g., device usage data) [478]. However, non-sensitive user data can become sensitive if linked to other data [311]. For example, smart switch usage in correlation with specific time periods of usage can infer sensitive user behavior data.
  • Device data: These are data related to IoT devices. Some user data stored in these devices is often associated with device data. However, device data is typically information that can be used to identify an IoT device within an IoT system [109, 180]. Examples of these devices include IP addresses, firmware versions, model number etc [52, 254].
2.4 IoT architecture
The deployment of IoT systems is highly flexible and depends on the specific requirements and types of IoT applications being utilized. For instance, when implementing IoT in a Smart Home (SH) environment, a middleware architecture or a gateway for internet connectivity might suffice. On the other hand, in an industrial IoT setup, a cloud-based architecture integrated with both fog and edge computing may be necessary [219]. Consequently, we outline the four primary IoT architectures, as highlighted in existing literature, applicable across diverse IoT applications.
2.4.1 Locally Hosted IoT system
In a traditional IoT system without cloud or Fog connectivity, a centralized software known as a middleware is essential to link the various IoT entities within the IoT ecosystem [502, 606]. This system offers a level of security when operating in a locally hosted environment, until the Gateway, an example of a middleware, permits external access to the Internet, making the locally hosted IoT system vulnerable to external threats [602, 124].
Middleware IoT system: It is an intermediary between an IoT system’s devices and application layers [606]. It provides communication and interaction management services among IoT devices, platforms, humans, and applications. Middleware is an essential part of an IoT ecosystem that enables the easy management of IoT components and provides scalability, security, and reliability for IoT applications by delivering timely updates to the latest, secure version [24].
A typical middleware example is the communication middleware that allows communication management between IoT devices, platforms, and humans using different sets of protocols and standards for data exchange [633]. Communication in IoT spans all entities that comprise the IoT ecosystem, as shown in Figure 2, and these types of communications are explained as follows.
  1. (i)
    Device-to-device (D2D): This is the autonomous communication between IoT devices without any human intervention, with centralized control for gathering, sharing, and multi-hop forwarding of information [71, 475].
  2. (ii)
    Device-to-cloud (D2C): This is the communication between the IoT device and the cloud service for storage purposes and management of the IoT system. The cloud can reliably store the IoT device data, with security, privacy, and uninterrupted availability [25, 475, 70].
  3. (iii)
    Device-to-application (D2A): This is interaction between IoT applications and IoT devices. This application can be installed on a user’s mobile device as an API manager or dashboard for controlling the IoT device or embedded into human life using content-centric, message technology, cloud technology, and sensor technology [475].
  4. (iv)
    Device-to-gateway (D2G): This is usually a secure communication between the local gateway and the IoT devices [475].
  5. (v)
    Device-to-Human (D2H) : This the communication between human and the IoT device. This communication passes user information to the IoT device, and vice versa [71].
Figure 2: The architecture of a locally-hosted IoT system showing communication types between user, devices, gateway(middleware), and cloud
2.4.2 Cloud-enabled IoT system
The IoT ecosystem recently depends greatly on cloud services through data storage, processing, and sharing support. This dependency is for adequate accommodation of ever-growing IoT data and service processing [447]. This allows data that cannot be stored locally on the IoT device due to limited storage capacity to be stored remotely in the cloud. The privacy of moving the data to the cloud and the privacy of the data at rest are essential [600]. Fog and edge computing are two common cloud-based approaches commonly used in IoT architecture, as shown in Figure 3.
2.4.3 Edge-enable IoT system
Edge computing system serves as a bridge between the physical and digital realms as shown in Figure 3. Devices at the edge gather real-world data, process it swiftly, and make it readily accessible to users via the internet [345]. In an edge-enabled IoT setup, the primary focus is on capturing sensor data and efficiently transmitting it to clients in a clear and understandable format [328]. Edge-enabled IoT offers scalability and enhances privacy in IoT systems by relocating applications from the cloud to edge servers, which are located closer to IoT devices.
2.4.4 Fog-enabled IoT system
Fog computing usually resides between cloud and edge computing, filtering out important data that needs to be processed in the cloud (see Figure 3). A fog-enabled IoT system involves edge computing, which brings control, storage, and networking closer to the IoT devices [375]. This setup enhances user experience without replacing cloud computing; it works alongside it. Fog computing significantly improves the quality of service (QoS) compared to users directly accessing the cloud [375]. Privacy requirements in a fog-enabled IoT system, as explained by Kinza et al. [431], revolve around content and context. Content privacy safeguards the data used for message delivery, while context privacy focuses on aspects of communication that could reveal the message’s content. Communication within a fog-based IoT system involves conveying content and context, necessitating the protection of both device and user data [431].
Figure 3: An IoT architecture showing the computing layer of the IoT system
2.5 Privacy, Privacy Types, Privacy threats, Privacy Preservation Techniques(PPTs), General Technology and Privacy Enhancing Technologies (PETs)
2.5.1 Privacy
It has come way back from the 14th century through the 18th, starting with one’s body and home protection, until it evolved into personal information access control and protection at the end of the 19th century [195]. Privacy is essentially the ability of individuals to control how much and under what circumstances they reveal their thoughts and actions to others. This concept is closely linked to one’s identity and is seen as a way to safeguard an individual’s autonomy and independence [245]. Privacy encompasses an individual’s right to control the collection, storage, and disclosure of their personal information. It serves to protect data, control how long data is retained, and regulate its use [35, 437]. When unauthorized exposure of this information occurs, it is known as a privacy breach. In the IoT context, a privacy breach refers to the disclosure of information about IoT users and devices to external entities, thereby posing a threat to individuals’ privacy [362].
2.5.2 Types of Privacy Protection
@@
Privacy in the IoT revolves around users and other entities interacting within IoT systems. Finn et al. [155] identify seven types of privacy in technology. These types include the privacy of the person, the privacy of personal behavior, the privacy of personal communication, the privacy of personal data, the privacy of location and space, the privacy of thoughts and feelings, and the privacy of associations [155]. However, not all of these privacy types are relevant to IoT. Therefore, based on existing literature, this study outlined the specific types of privacy in IoT systems.
@@
  • Device Identity privacy: To prevent unauthorized device tracking and smart device information disclosure, such as firmware version, operating system, MAC address, etc. [299].
@@
  • Location privacy: To protect the user and device location information. This could prevent threats such as profiling [155, 299].
@@
  • Group privacy: To protect the privacy of a group of users [155]. For example, a family member of a Smart Home (SH) owner or department member in an Industrial IoT (IIoT) environment who has access to operate an automated system in an industry.
2.5.3 Privacy Threats in IoT
@@
Depending on the threat actor, some privacy threats can be malicious, while some can be non-malicious. Privacy threats, such as subject identification, linkage, data leakage, inventory attacks, eavesdropping, life cycle transition, privacy violations during presentation, and data tampering, can be malicious due to the threat actor involved [362, 645, 600, 437]. In contrast, privacy threats such as Localization and tracking, and profiling are not usually malicious when they involve actors such as service providers, security agents, and third-party providers [362, 645]. These privacy threats could affect either IoT users or devices, or both
2.5.4 Privacy Preservation Techniques (PPT)
@@
PPTs are methods or processes implemented in IoT to achieve privacy goals [459]. These privacy goals vary depending on the techniques and the IoT application involved. These privacy goals include data masking, pseudonymization, data perturbation, multi-path routing, and access controls. To achieve this goal by PPTs, the right technological building block needs to be developed, which is referred to as privacy-enhancing technologies (PET) [431, 6, 509, 15]. PET in IoT may be challenging due to the constraints of IoT systems, such as limited memory capacity and processing power. However, implementing these PETs in a lightweight manner will make them suitable for IoT systems [459].
2.6 Privacy-Enhancing Technologies (PET)
@@
These are technologies used to achieve PPTs in an IoT system. In IoT, PPTs have evolved, employing diverse PETs (such as differential privacy, homomorphic encryption, and anonymization, to mention a few) to provide various privacy preservation solutions [264, 439, 87, 194, 272, 422, 246].
2.7 General Technology used in privacy Privacy Preservation Approach in IoT
@@
The Internet of Things (IoT) utilizes various privacy preservation approaches to achieve specific privacy goals and types. These approaches often leverage PETs developed through general technological approaches like blockchain [132], machine learning/deep learning [642], cryptography [36], informed consent [170], quantum-safe algorithms [426], mathematical models [366], and privacy-based infrastructure.
The overall privacy preservation taxonomy and the involved entities are represented in Figure 4
3 Comparison of this paper with other surveys and review papers
In this section, we comprehensively analyzed areas previously explored by surveys focusing on PPTs within IoT. These areas were classified based on the technology implemented, the goals of PPTs, the specific IoT applications in which these techniques were applied, and the types of privacy they aimed to protect. The comparative insights shown in Table 1 highlight the uniqueness of this scoping review. This study is the first to conduct a comprehensive scoping review of PPT in an IoT system, providing a detailed understanding of the technologies used to build PETs, IoT applications, privacy goals, the implementation of PPT across different layers of IoT computing, and the privacy types associated with these techniques. We present an explanation of similar reviews related to this study as follows:
Table 1: Comparison of Our work to Prior Reviews on Privacy Preservation in IoT